OpenStudy (anonymous):

How do you multiply these? 5 1/5 x 4 1/4 =

OpenStudy (anonymous):

Please explain in steps.

myininaya (myininaya):

Write the mix fractions as improper fractions like so: \[Q \frac{R}{D} =\frac{D \cdot Q+R}{D}\]

myininaya (myininaya):

\[\frac{5 \cdot 5+1}{5} \cdot \frac{4 \cdot 4+1}{4}=\frac{25+1}{5} \cdot \frac{16+1}{4}=\frac{26}{5} \cdot \frac{17}{4}\]

OpenStudy (anonymous):

3 4/5 x 2 2/8 =

myininaya (myininaya):

\[\frac{26}{5} \cdot \frac{17}{4}=\frac{26}{4} \cdot \frac{17}{5}=\frac{13}{2} \cdot \frac{17}{5}=\frac{ 221}{10}\]

OpenStudy (anonymous):

5 1/5= 5.2 4 1/4= 4.25 5.2 X 4.25= 22.1

myininaya (myininaya):

I will be happy to check your answer. Look what I did above and try to do that here or you can do what michu is doing which is writing both as decimals and then multiplying. (his second reply is right :))

OpenStudy (anonymous):

Just change both of the fraction to improper fraction and the divide it by the denominator and lastly you would multiply it.